# Large Language Models Can Learn Temporal Reasoning

### v5

@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have emerged with some reasoning capabilities  (Huang and Chang, 2023). However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems. In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Vashishtha et al., 2020; Qin et al., 2021).
@@
Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2022). Similarly,  (Wei et al., 2023) designs several specific types of prompts as prompt tuning.  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Given a reliable TG, the key challenges of teaching TR to LLMs include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? Prior work  (Lewis et al., 2021) shows that explicitly introducing knowledge into context enhances the performance of LLMs. In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., temporal relation and time gap between the timestamps, and the relative order of the gaps). (2) How can one teach LLM to perform deliberate reasoning? Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b). In this paper, we adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. We further improve the model performance with graph data augmentation, which mitigates the data deficiency in TR tasks.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFT on reliable CoTs brings better reasoning performance than that on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). Since asking humans to create the CoT data is not scalable, we use LLM to replace humans. This task is non-trivial for reasoning over knowledge graphs  (Saparov and He, 2023). In this section, we propose a bootstrapping pipeline, i.e., given a query, using LLM to generate several CoTs and selecting them as training data with a weighted sampling strategy. Compared with the conventional Best-of-N sampling, our proposal allows more training data diversity.
@@
$\log P\left(a\mid q^{\dagger},c_{k}\right)=\frac{1}{|a|}\sum_{l}^{|a|}\log P\left(t_{l}\mid q^{\dagger},t_{<l}\right)$ (4)
@@ 3.2.2 Graph Data Augmentation
Compared with other tasks, reasoning suffers more from data insufficiency since more information (such as evidence, arguments, and logics) are involved in the intermediate steps  (Huang and Chang, 2023). To address this issue, we propose several graph data augmentation strategies (Figure 3).
@@ 4.1 Experimental Setup
We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al. (2023b), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023). Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models. We evaluated their few-shot ICL performance on the test set. To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
@@ 5 Related Work
Language-based Temporal Reasoning. Recently, language-based TR has gained substantial research focus  (Liu et al., 2023, 2024b; Chen et al., 2024a, b; Wang et al., 2024; Cheng et al., 2024a; Xia et al., 2024). The vision here is to help LMs understand temporal concepts and logic such that they can perform more complicate tasks. Existing methods mainly solve this problem with time-aware language modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose specific pre-training/fine-tuning strategies for robust TR. On the other hand,  (Ning et al., 2019; Zhou et al., 2020, 2021; Yang et al., 2020, 2023a) design auxiliary objectives to introduce external temporal knowledge. Although these methods made some progress, representation learning for the underlying structure and logic of TR are either ignored or not explicitly involved.
@@
Reasoning towards LMs. Although LLM has exhibited some emergent behaviors (Zhang et al., 2024a; Lai et al., 2024a, b; Lyu et al., 2024), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is. Existing methods that try to elicit or enhance reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use in-context learning where some demonstrations including intermediate reasoning steps are provided in prompt.  (Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with the intermediate thinking process before producing final answers. On the other hand,  (Schick et al., 2023; Kynoch et al., 2023; Xu et al., 2023) combine LMs with domain-specific external tools, empowering the model to perform more complex tasks that require reasoning and interactions with environment.
Reasoning over Knowledge Graphs. Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Sun et al., 2018; Yang and Song, 2020; Huang et al., 2023; Wan et al., 2024; Li et al., 2024). Related work includes symbolic reasoning over temporal KGs  (Yang et al., 2022; Xiong et al., 2023, 2024), and language-based reasoning over static KGs  (Cheng et al., 2024b; Zhang et al., 2024b; Liu et al., 2024a; Zhao et al., 2024; Xu et al., 2024; Wei et al., 2024). In this paper, we build the connection between language-based and symbolic-based TR. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to graph reasoning.

### v6

@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery. Recently, large language models (LLMs)  (Ouyang et al., 2022; Achiam et al., 2023; Touvron et al., 2023a) have emerged with some reasoning capabilities  (Huang and Chang, 2022). However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems. In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Vashishtha et al., 2020; Qin et al., 2021).
@@
Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2022). Similarly,  (Wei et al., 2023) designs several specific types of prompts as prompt tuning.  (Li et al., 2023; Yuan et al., 2024) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Given a reliable TG, the key challenges of teaching TR to LLMs include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? Prior work  (Lewis et al., 2020) shows that explicitly introducing knowledge into context enhances the performance of LLMs. In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., temporal relation and time gap between the timestamps, and the relative order of the gaps). (2) How can one teach LLM to perform deliberate reasoning? Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b). In this paper, we adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. We further improve the model performance with graph data augmentation, which mitigates the data deficiency in TR tasks.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFT on reliable CoTs brings better reasoning performance than that on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2022). Since asking humans to create the CoT data is not scalable, we use LLM to replace humans. This task is non-trivial for reasoning over knowledge graphs  (Saparov and He, 2022). In this section, we propose a bootstrapping pipeline, i.e., given a query, using LLM to generate several CoTs and selecting them as training data with a weighted sampling strategy. Compared with the conventional Best-of-N sampling, our proposal allows more training data diversity.
@@
$\log P\left(a\mid q^{\dagger},c_{k}\right)=\frac{1}{|a|}\sum_{l=0}^{|a|}\log P\left(t_{l}\mid q^{\dagger},t_{<l}\right)$ (4)
@@ 3.2.2 Graph Data Augmentation
Compared with other tasks, reasoning suffers more from data insufficiency since more information (such as evidence, arguments, and logics) are involved in the intermediate steps  (Huang and Chang, 2022). To address this issue, we propose several graph data augmentation strategies (Figure 3).
@@ 4.1 Experimental Setup
We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al. (2023b), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023). Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models. We evaluated their few-shot ICL performance on the test set. To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
@@ 5 Related Work
Language-based Temporal Reasoning. Recently, language-based TR has gained substantial research focus  (Liu et al., 2023, 2024b; Chen et al., 2024a, b; Wang et al., 2024; Jiayang et al., 2024; Xia et al., 2024). The vision here is to help LMs understand temporal concepts and logic such that they can perform more complicate tasks. Existing methods mainly solve this problem with time-aware language modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose specific pre-training/fine-tuning strategies for robust TR. On the other hand,  (Ning et al., 2019; Zhou et al., 2020a, b; Yang et al., 2020, 2023a) design auxiliary objectives to introduce external temporal knowledge. Although these methods made some progress, representation learning for the underlying structure and logic of TR are either ignored or not explicitly involved.
@@
Reasoning towards LMs. Although LLM has exhibited some emergent behaviors (Zhang et al., 2024a; Lai et al., 2024a, b; Lyu et al., 2024), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is. Existing methods that try to elicit or enhance reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2024) use in-context learning where some demonstrations including intermediate reasoning steps are provided in prompt.  (Nye et al., 2021; Wang et al., 2023; Ho et al., 2022) fine-tune LMs with the intermediate thinking process before producing final answers. On the other hand,  (Schick et al., 2024; Kynoch et al., 2023; Xu et al., 2023) combine LMs with domain-specific external tools, empowering the model to perform more complex tasks that require reasoning and interactions with environment.
Reasoning over Knowledge Graphs. Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Huang et al., 2023; Wan et al., 2024; Li et al., 2024). Related work includes symbolic reasoning over temporal KGs  (Yang et al., 2022; Xiong et al., 2023, 2024), and language-based reasoning over static KGs  (Cheng et al., 2024; Zhang et al., 2024b; Liu et al., 2024a; Zhao et al., 2024; Xu et al., 2024; Wei et al., 2024). In this paper, we build the connection between language-based and symbolic-based TR. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2024; Gao et al., 2024), which limit their application to certain tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to graph reasoning.